OpenStudy (anonymous):

If g(x) = 2(x − 4), find the value of x if g(x) = 20. 32 12 14 10

OpenStudy (anonymous):

14

OpenStudy (anonymous):

can u explai wanna learn '

OpenStudy (anonymous):

plug 20 in for g(x). so it will be 20=2(x-4). distribute the 2 to both the x and the -4 which gets you 20=2x-8. Then add 8 to both sides which gets 28=2x. finally divide both sides by 2 which =14

OpenStudy (anonymous):

thnk u sir

OpenStudy (anonymous):

you are welcome!!

OpenStudy (anonymous):

can u help me with another one

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

If f(x) = x + 2, find f(3). 1 6 5 −1

OpenStudy (anonymous):

plug 3 in for x so you get f(x)=3+2 which =5

OpenStudy (anonymous):

ure a life savor thnks

OpenStudy (anonymous):

you are so welcome!!

OpenStudy (anonymous):

If g(x) = x2 + 2, find g(3).

OpenStudy (anonymous):

i got 11

OpenStudy (anonymous):

plug 3 in for x which gets you g(x)=(3*2)+2 which gets 6+2

OpenStudy (anonymous):

which equals 8

OpenStudy (anonymous):

i ment 8 lol u should be a totur

OpenStudy (anonymous):

im here for you anytime.

OpenStudy (anonymous):

If h(x) = 2x + 1, find the value of x if h(x) = 11.

OpenStudy (anonymous):

plug 11 in for h(x) which gets you 11=2x+1. Then subtract 1 from both sides which gets you 10=2x. finally divide both sides by 2 which gets x=5

OpenStudy (anonymous):

thts wht i go to lol nice are u good in biology

OpenStudy (anonymous):

yes i am!!

OpenStudy (anonymous):

ask me anything!

OpenStudy (anonymous):

■Detailed descriptions of what happens to the carbon atom: ■during photosynthesis (the light-dependent reactions and the light-independent reactions) ■transitioning between photosynthesis and respiration ■during cellular respiration or fermentation ■An explanation of how the carbon atom is transferred between molecules. Please include descriptions of the molecules involved in each of the processes above and be sure to mention the movement of electrons and energy as the carbon atom moves from one reaction to another.

OpenStudy (anonymous):

its an essay cant get my mind around it only 3 paragrapths if u can help tht would be great

OpenStudy (anonymous):

let me see what i can do!

OpenStudy (anonymous):

ure the best no joke

OpenStudy (anonymous):

Step 1: Carbon Fixation. This phase begins when a carbon dioxide molecule is attached to a 5 carbon sugar, ribulose biphosphate (RuBP). This reaction is catalyzed by the enzyme RuBP carboxylase (rubisco) one of the most abundant proteins on earth. The products of this reaction is an unstable 6 carbon compound that immediately splits into 2 molecules of 3-phosphoglycerate. For every 3 molecules of carbon dioxide that enter the cycle via rubisco, 3 RuBP molecules are carboxylated forming 6 molecules of 3-phosphoglycerate. Step 2: Reduction. This endergonic reduction phase is a 2 step process that couples ATP hydrolysis with the reduction of 3-phosphoglycerate to glyceraldehyde phosphate. An enzyme phosphorylates (adds a phosphate) 3-phosphoglycerate by transferring a phosphate from the ATP. The product is 1-3-bisphosphoglycerate. Electrons from the NADPH reduce the carboxyl group of the 1-3-bisphosphoglycerate to the aldehyde group of glyceraldehyde-3-phosphate. For every three carbon dioxide molecules that enter the Calvin cycle, 6 glyceraldehyde-3-phosphates are produced, only one can be counted as a net gain. The other 5 are used to regenerate 3 molecules of RuBP. Step 3: Regeneration of RuBP. A complex series of reactions rearranges the carbon skeletons of 5 glyceraldehyde-3-phosphate molecules into 3 RuBP molecules. These reactions require 3 ATP molecules. RuBP is thus regenerated to begin the cycle again.

OpenStudy (anonymous):

how would i sum tht up in like 3 pargrapths or 2

OpenStudy (anonymous):

A case could exist where a covalent bond forms between carbons of two different molecules, A and B, so that they become one molecule. If a different bond then broke in what had been Molecule B so that it dissociated, leaving part of the initial molecule bound to A, the net result would be the "transfer" of one or more carbon atoms from molecule B to molecule A. But in general, molecules don't "trade" or "transfer" whole atoms. Particularly in biologic systems, where energy barriers involved in bond making and breaking events are, in general, high.

OpenStudy (anonymous):

this website should help you a lot! http://www.cliffsnotes.com/study_guide/Process-of-Photosynthesis.topicArticleId-8741,articleId-8599.html

OpenStudy (anonymous):

ure the besttttttttttttttttttttttt it would be rude to ask for anymore lol

OpenStudy (anonymous):

no not at all! do you have any other questions?

OpenStudy (anonymous):

like 15 questions if ure up for it lol

OpenStudy (anonymous):

of coarse! lay them on me!